### v1

# Recipes for Pre-training LLMs with MXFP8

Abstract
Precision scaling — using fewer bits to represent model parameters and related tensors during pre-training — has emerged as a compelling technique for improving GPU efficiency without sacrificing accuracy. Microscaling (MX) formats [1] in NVIDIA’s latest Blackwell GPUs represent a major leap in enabling this precision scaling aspect. These formats combine narrow floating-point data types with per-block scaling factors, offering a fine-grained approach to quantizing tensors.
Although MX-formats offer the promise of improved numeric stability compared to other reduced-precision representations, in practice they must be used carefully in order to successfully converge an LLM on a multi-trillion token dataset. In this paper, we show that the rounding mode suggested in OCP [1] specification can lead to divergence when pre-training an LLM. We show an improved rounding mode, which uses round-to-infinity to compute scaling factors, enables successful pre-training in MXFP8 for an 8B model on 15T tokens.
1 Introduction
Scaling the number of parameters of deep learning models, especially large generative models, has intensified the need for more efficient compute and memory solutions. Precision scaling — reducing number of bits to represent a data type — is a successful strategy to improve performance while lowering memory requirements. However, precision scaling without sacrificing model accuracy remains a key challenge.
Microscaling (MX) formats, developed in the Open Compute Project (OCP) [1], aims to offer a standardized approach to low-precision representation. By combining narrow floating-point types with fine-grained scaling factors, MX-formats aim to deliver a balance of dynamic range, precision, and hardware efficiency. OCP [1] and research works like [2, 3, 4] have shown a fine-grained scaling approach to be an effective strategy compared to granular (e.g., per-tensor) scaling approaches, especially in sub-8-bit pre-training regimes.
NVIDIA’s latest GPU generation, Blackwell [5], adds native MX data type support to Tensor Cores. Blackwell supports MX data for 8-bit (MXFP8), 6-bit (MXFP6) and 4-bit (MXFP4) floating-point types. This paper primarily focuses on MXFP8 and provides an in-depth analysis of pre-training large language models (LLMs). It details the conversion process from high precision formats to the quantized MXFP8 format, including rounding modes for data and scaling factors. These insights aim to guide the efficient and accurate pre-training of LLMs using MXFP8.
Notably, we observe that the rounding mode suggested in OCP [1] for scale factor computation causes MXFP8 pre-training to differ from a BF16 baseline. We propose a modification to this scale factor rounding where we round-to-infinity instead of rounding-down, and demonstrate that this adjustment enables MXFP8 pre-training to match BF16 pre-training accuracy.
The primary contributions of this paper are as follows:
  • A methodical study of MX quantization for LLM pre-training where we investigate the following design choices: data formats, quantization schemes, storage, the selection of tensors to quantize, and examining how these choices affect accuracy. Specifically, we identify E4M3 to be the MXFP8 data type that best helps maintain pre-training accuracy.
  • We find that the rounding mode suggested in the OCP specification can lead to divergence during pre-training. Building on the analysis, we propose a rounding mode for scale factors that lets MXFP8 trained models to match BF16 model accuracy. Specifically, we describe a recipe to pretrain LLMs with MXFP8 and show experiments where an 8 billion parameter dense LLM trained on 15 trillion tokens matches a BF16 trained model and a per-tensor scaled FP8 [6] model with first and last layer in BF16.
For each design choice, we present ablation studies on small models (e.g. 100s of millions of parameters) pretrained on short (e.g. 100s of billions) token horizons. We then apply the findings to large multi-billion parameter models trained on trillion token horizons. This makes our conclusions robust for large-scale foundation model pre-training.
2 Microscaling format support in NVIDIA Blackwell
Figure 1: A single MXFP block (in green box) and interpretation of MXFP format.
Background: An MX-format is specified by a block size K, a shared scaling factor per block, X, and the data-type of elements in the block. A block is a contiguous set of K elements along an axis in a tensor. K = 32 for all MX types on Blackwell. The data-type of X is UE8M0, an unsigned exponent-only value that encodes either NaN or any power-of-two [^1] Power-of-two is a number of the form $2^{n}$ where n is a positive integer, negative integer, or zero. value in range $2^{-127}$ to $2^{127}$.
Given K input elements in a source-format, $V_{i}$ (typically FP32); $1\leq i\leq K$, conversion to MX-format consists of computing X and $Q_{i}$ such that $Q_{i}\times$X is the decoded value corresponding to $V_{i}$. X and $Q_{i}$ are stored in memory instead of $V_{i}$. This is depicted in Fig. 1. A tensor in source-format is sub-divided into blocks of K elements and converted into MX-format to be stored in memory and/or processed by math units in hardware. Tensor Cores in Blackwell consume X and $Q_{i}$ to compute the dot product of two MX-formatted blocks. If the accumulated output of a dot-product operation in Tensor Cores is FP32, then this output is thereafter quantized to MX-format if a subsequent operation consuming the output needs it in that format. The conversion process is, $Q_{i}=\verb+Quantize_to_fp8+(V_{i}/2^{X})$. Section 3 describes the conversion details. Fine-grained scaling helps each MX block independently align to the needed range of input values before quantization.
Table 1 shows the MX-formats supported in Blackwell. The data type column uses the convention ExMy to denote x-bit for floating-point exponent and y-bit for mantissa. For a fixed bit-width, floating-point numbers trade-off exponent width with mantissa width. Thus, MXFP8 E4M3 is a 8-bit data type with with 1 sign bit, 4-bit for FP exponent and 3-bit of mantissa. Similarly, E5M2 is an 8-bit type with 5-bit for exponent and 2-bit for mantissa. Compared to E4M3, E5M2 can represent a larger dynamic range, but with less precision. E5M2 is the only type that follows IEEE 754 conventions [7] for representation of special values. For all other data types, dynamic range is extended by not representing Infinity and NaN (E4M3 has only one bit-pattern for NaN [6]). More bits in the exponent field translates to a larger range while more bits in the mantissa field translates to more precision within a given range. Every floating-point type has a dynamic range that it can represent — we denote this range in terms of binades which is the log2-ratio of the maximum to the minimum finite representable in that format.
Table 1: MX-format support in Blackwell
Format Data Max. Min. Binades Relative speed
type normal subnorm. vs BF16 (Tensor
Core math)
MXFP8 E4M3 1.75 * $2^{8}$ $2^{-9}$ 17.8 2x
E5M2 1.75 * $2^{15}$ $2^{-16}$ 31.8 2x
MXFP6 E2M3 1.875 * $2^{2}$ $2^{-3}$ 5.9 2x
E3M2 1.75 * $2^{4}$ $2^{-3}$ 8.8 2x
MXFP4 E2M1 1.5 * $2^{2}$ $2^{-1}$ 3.6 4x
3 Pre-training with MXFP8
Model weights, activations, and gradient tensors are quantized from FP32 into MXFP8 during the forward and backward passes in training. MXFP8 tensors are then stored and operated on by the hardware. We first describe the conversion process, Quantize_to_fp8$(V_{i}/2^{X})$, starting with scale computation and following on to the quantization procedure for each element. The quantization process described below is the same across all MX-formats (E4M3, E5M2 , E2M3, E3M2, and E2M1), with only the MX data type differing.
3.1 Conversion from FP32 to MXFP8
Computing X: Typically, most of the values within each block in a tensor will be out of the representable range for the target MX format, both underflowing the minimum and overflowing the maximum representable number. To address this, a scale factor is multiplied with all values within a block, to shift the maximum number of values to be within the representable range.
The scale factor, X, is computed based on the absolute maximum value (amax) among the 32 high-precision input values, i.e. amax = $\max\|V_{i}\|;1\leq i\leq 32$. The goal is to map this amax in the input-format to be the largest representable value in the desired MX-format. Special care is taken if some or all the elements in the input are Infinity and/or not-a-number (NaN). If amax in a block is 0, then X is set to -127, thus, interpreted value of X is $2^{-127}$. All $Q_{i}$’s in this case are set to 0.
When X is not Inf, NaN or 0, the OCP [1] specification suggests X be set to the largest power-of-two less than or equal to amax divided by the largest power-of-two representable in the MX-format type. For example, with the E4M3 type, $X+=\frac{\text{floor}(\log_{2}(\texttt{amax}))}{\text{floor}(\log_{2}(448))}$, since 448 is the largest magnitude in E4M3. Thus, the OCP specification ignores the floating-point significand of this ratio. [^2] Assume amax and max representable in MX-format (destamax) are normal floating-point numbers of the form amax = $2^{A}$ 1.$M_{a}$ and destmax = $2^{E}$ 1.$M_{e}$ (where mantissa $M_{a}$ and $M_{e}$ lie between 0 (inclusive) and 1. OCP rounding scheme results in scale exponent = A-E (ignoring the case where min(A-E) is clamped to -127)
We observe accuracy degradation when following the OCP specification. Fig. 2 show training loss curves for a 843 million parameter transformer model trained under two token horizons: 300 billion and 1 trillion tokens. Two different configurations are studied:
  • (cfg1) E4M3 for all tensors (weights (W), activations (A) and gradients (G)), and
  • (cfg2) E4M3 for W and A tensors, E5M2 for G tensors.
The E5M2 format has $\sim$1.8x more binades than E4M3 and since gradients typically have larger dynamic range, [6] advocated the use of E5M2 with per-tensor scaling method. This rationale underpins the study of E5M2 gradients. In both cfg1 and cfg2, scale factor computation using OCP method leads to training divergence (Fig. 2(a)) or a widening loss gap relative to BF16 (Fig. 2(b)).
(a) Validation perplexity curves for 843M parameter model trained on 300B tokens.
(b) Validation perplexity curves for 843M parameter model trained on 1T tokens.
Figure 2: Comparing scale factor rounding modes suggested in the OCP specification and our work. E4M3 and E5M2 are MXFP8 formats; notation E4M3(W) denotes E4M3 data type for Weights. Similarly, E4M3(A) and E4M3 (G) refer to activations and gradients, respectively.
Algorithm 1 outlines our method for computing the scale factor. The key change is to round-up towards positive infinity the exponent of ratio of amax and max representable in MX-format, destmax (while saturating to UE8M0 max/min limits). This is in contrast to the OCP scheme which effectively suggests to round-down the scale value. Since a high-precision value, $V_{i}$, is scaled by the scale factor $2^{X}$, rounding-up the denominator of the fraction ($V_{i}/2^{X}$) has a tendency to map amax below destmax. In contrast, the OCP method has a tendency to map amax above destmax (which subsequently needs to be clamped so that it becomes representable). We hypothesize that this saturating effect with OCP rounding method affects model accuracy. A more detailed and accurate workflow to compute and round the value of X is described in Appendix Sec. A.1.
Algorithm 1 Our proposal for computing scale factor, X (simplified method)
$X_{\text{float}}\leftarrow$ amax/destmax $\triangleright$ destmax is max representable in MX-format
$\text{expX}_{\text{float}}\leftarrow\log_{2}(X_{\text{float}})$triangleright extract the exponent of X (de-biased form)
$\text{expX}_{\text{int}}\leftarrow\text{ceil}(\text{expX}_{\text{float}})$triangleright round-up
$X\leftarrow\text{clamp}(\text{expX}_{\text{int}},-127,127)$triangleright clamp to min/max E8M0 representable
$X\leftarrow X+127$triangleright add bias
return $X$triangleright store X
Fig. 2 shows that MXFP8 with E4M3 for gradients (blue curve) and MXFP8 with E5M2 for gradients (purple curve) with our proposed rounding scheme both overlap with a reference BF16 loss curve across both 300B and 1T token horizons. Later, in section 3.2 we discuss that E4M3 is in fact a better choice for pre-training LLMs versus E5M2 with MX style fine-grained scaling.
Quantizing FP32 values to MX type: Once X is computed, $V_{i}$ is scaled by $2^{X}$ and the resulting value is quantized to a FP8-representable number. This is the Quantize_to_fp8() function. Round-to-nearest-ties-to-even (RN) rounding is used during this quantization step. The conversion process is saturating, i.e. if after rounding the resulting value exceeds FP8 max or is less than FP8 min value, then the result is clamped to respective max or min value.
A practical instance of this conversion process in low-precision LLM pre-training arises when matrix multiplication and accumulation (MMA) outputs, typically stored in FP32, must be mapped to MXFP8. In this case 8-bit quantized values are stored in memory, thus saving write bandwidth and storage capacity when compared with storing FP32 values. Subsequent model operation then reads MXFP8 values, saving read bandwidth compared to loading FP32 values. Further, since Tensor Cores can process MX-formatted inputs, the MMA operation in lower precision consumes less energy and operates at higher throughput.
3.2 E4M3 data type for weights, activations and gradients
(a) Validation perplexity curves for 843M parameter model trained on 1T tokens.
(b) Validation perplexity curves for 8B parameter model trained on 1T tokens.
Figure 3: Pre-training loss curves comparing E4M3 and E5M2 when used across different tensor types: weights (W), activation (A) and gradients (G). The inset shows a zoomed-in view of the loss at the end of training.
FP8 has two variants in Blackwell 1: E4M3 and E5M2. Our experiments show the following:
  • Better convergence with E4M3 instead of E5M2 for weights and activations during training. Fig. 3(a) shows the loss behavior comparing E4M3 and E5M2 for the same 843M model used in Fig. 2(b). The purple curve has E5M2 for activations and the blue curve has E5M2 for weights while all other tensors are in E4M3. Both these curves show worse loss convergence compared to using E4M3 for all tensors (orange curve) or using E5M2 for gradients (yellow curve).
  • E4M3 for gradients maintains training loss parity with BF16 pre-training, especially for models with 2 billion or more parameters. Fig. 3(b) shows the loss behavior comparing E4M3 and E5M2 for the gradient tensor on an 8 billion parameter LLM trained on 1 trillion tokens: using E4M3 gradients (orange) has lower loss than using E5M2 gradients (yellow). This gap increases as the model is trained on more tokens. This change in behavior with increasing model parameter counts underscores the importance of examining the numerical properties of formats across a broad range of model sizes.
Previously, per-tensor scaled FP8 studies in [8, 6, 9] and coarse-grained block-scaled FP8 studies in [10] used E5M2 for gradient tensors in lieu of E4M3. With fine-grained scaling, the dynamic range requirements at a 32-element block size is sufficiently captured by 17.8 binades in E4M3 type. Once the range requirements are met, precision (or sampling) becomes important, and E4M3 having 8 samples per binade is better than E5M2 having only 4. Thus, for our MXFP8 pre-training recipe, we quantize all three tensor types – weights, activations, and gradients – with the E4M3 data type.
Figure 4: Top: Transformer layers quantized to MXFP8 inside a single transformer block. Bottom: Training workflow for a single layer during FPROP, DGRAD and WGRAD.
Layers to instantiate in MXFP8 and training workflow: We use language-based transformer models for all studies in this paper. Our future work will study these recipes on speech and vision models. Based on our studies, our guideline is to quantize QKV, PROJ and FFN Up- and Down-projection layers to MXFP8 across all transformer blocks in the model. The Batch-Matrix Multiplications (BMM1, the query-key dot product and BMM2, the attention score-value product) in the self-attention layer, along with operations like Softmax, Act-func and residual-add are in high-precision. We found this to be the safest option for maintaining accuracy parity with BF16 pre-training. This is illustrated in Fig. 4. The input embedding layer and the final output-projection layer are also in BF16 or FP16. All studies in this paper use this guideline.
During training, with MXFP quantization, the training framework has to keep two copies of each of the tensors – weights, activations and gradients: each copy is quantized along the axes of dot-product reduction (row and column). Fig. 4 shows how each tensor is used in forward (FPROP), weight-gradient (WGRAD) and activation-gradient (DGRAD) computation during the training loop. Since each tensor is used in non-transposed and transposed form, quantization needs to occur along two separate axes (row and column).
Summary so far: We introduce a new rounding scheme for the MX scale factor that addresses the divergence caused by the OCP-based approach and achieves loss parity with BF16 on an 843M-parameter model up to 1T tokens. Additionally, adopting E4M3 and our scale factor calculation method in Alg. 1 enables scaling to an 8B-parameter model trained on 15T tokens — representing the largest LLM pre-training at this scale with MXFP formats, to the best of our knowledge. Appendix (section A.4) shows that our recipe also holds for a 16B-parameter mixture-of-experts model.
3.3 Results with MXFP8 pre-training a LLM on 15T tokens
Figure 5: Pre-training a 8B LLM on 15T tokens. Top: Training behavior of BF16 vs MXFP8. Bottom: Comparing BF16, FP8 and MXFP8’s downstream task scores on MMLU and a set of 9 reasoning tasks. MXFP8 numerics uses our proposed rounding method and E4M3 for all quantized tensors.
We pre-train an 8B parameter Nemotron model [11] using Megatron-LM [12]. The model has 32 transformer blocks, 32 attention heads, hidden size is 4096, GQA group size is 8, KV-channels is 128, sequence length during pre-training is 8192. It is trained on 15T tokens with a batch size of 768. Initial learning rate is 6e-4 that cosine decays to 6e-6. A phased data-blending approach is used to train the model: in the first phase, a data mixture that promotes diversity in data is used and in the second phase high-quality datasets (e.g., Wikipedia) are used. We switch to the second phase at the 60% point of training. This blend style has also been used in other large scale pre-training setups [8]. The model is pre-trained on 3072 Hopper GPUs (since hardware with MX support was unavailable during much of the experimentation duration).
We use an emulation of MX-formats on Hopper GPUs: tensors that feed a MMA operation are first quantized into the MX-format and then cast back to BF16 before the BF16 MMA operation. The training workflow depicted in Fig. 4 is implemented in Megatron-LM. We validated the numerical fidelity of our emulation package by comparing with a 2B parameter LLM pre-training run on Blackwell using actual MXFP8 format and confirmed that they were identical.
Fig. 5 shows training loss behavior and task-level accuracy for the 8B pre-trained model. We report evaluation scores on two sets of downstream tasks: (1) 5-shot score on MMLU [13] and (2) Averaged 1-shot score across 9 general reasoning benchmarks: ARC-Challenge and ARC-Easy [14], Race [15], PIQA [16], Winogrande [17], Hellaswag [18], OpenBookQA [19], Social IQA [20] and Commonsense QA [21]. We observe the following:
  1. Validation perplexity of the model when pre-trained with MXFP8 matches pre-training with BF16 (left plot in Fig. 5). There is less than 0.50% difference between MXFP8 and BF16 validation perplexity values throughout the pre-training run.
  2. The middle and right plots in Fig. 5 show evaluation scores on two sets of downstream tasks. Again, scores for the MXFP8 trained model match with those of the BF16 trained model. This makes MXFP8 a viable candidate for pre-training LLMs.
MXFP8 versus FP8: In addition to MXFP8 and BF16, Fig. 5 also shows task-level scores for the same model trained with traditional FP8 precision. FP8 recipe uses software-emulated block scaling [6] where an entire tensor is scaled so that maximum number of tensor values fall into the representable range in the quantized format. We follow the guidelines suggested in [8] for FP8 pre-training setup: the first and last transformer block in the model are kept in BF16 while linear layers of the remaining blocks are quantized to FP8.  [8] found this choice appropriate for pre-training 8B and 56B parameter LLMs on 20T tokens. Keeping some layers in BF16 affects end-to-end speedup and also complicates pre-training — a choice has to be made on which layers to leave in higher-precision. We observe that MXFP8 matches FP8 accuracy on these two sets of tasks without requiring any BF16 layers.
MXFP8 versus blockwise-FP8: Further, some works like Deepseek-V3 [10] report the need to scale to a smaller block size when using FP8. In this setup, certain tensors require vector 1x128 scaling and certain tensors require per-block (e.g. 128x128) software scaling. This complicates GEMM kernel design. Native support for MXFP8 simplifies this — fine-grained scaling provides better numerical robustness and avoids any tradeoff between smaller block sizes and hardware speed.
In summary, we find MXFP8 maintains accuracy compared to BF16 or FP8 pre-trained models. On GB200 Blackwell systems, MXFP8 has 2$\times$ higher throughput than BF16 making end-to-end MXFP8 pre-training faster than BF16 pre-training. We also find the MXFP8 recipe to be simpler to use when compared to FP8 (all layers can be quantized and scaling is handled in the hardware) while allowing for equal or better throughput.
4 Related work
Low-precision training and inference is a widely studied topic in deep learning literature. While significant progress has been made in low-precision inference [22, 23, 24, 25], there are relatively fewer studies demonstrating low-precision techniques for pre-training LLMs, especially large-scale pre-training on large token horizon. Our work primarily focuses on 8-bit pre-training and prior work on related low-precision LLM pre-training can be grouped into the following two categories:
  1. LLM pre-training using FP8 formats: [6] proposes an FP8 binary interchange format, consisting of E4M3 and E5M2 encodings, and a per-tensor scaling approach — an entire tensor is scaled to capture the maximum number of tensor values in the representable range in FP8[26] discusses FP8 pre-training challenges and proposes model-level modifications to train a 7B parameter model. Recently, per-tensor scaled FP8 was used to train the Nemotron-H family of LLMs [8] and the Llama-4 family of models also used FP8  [27]. Instead of per-tensor scaling, DeepSeek-V3 family of models [10] use block-scaled FP8, this helps to better capture outliers and minimize quantization errors.
  2. Pre-training using MXFP formats: [2] presents empirical data on pre-training models with MXFP formats. They show cast-only inference results for MXFP8 and pre-training results for MXFP6 and MXFP4-weights. [28] studies MXFP4 backward-pass quantization and [29] investigates MXFP4 weight quantization on relatively small token horizon. All these studies are based on the scaling factor computation method described in [1] which we show to be ineffective at large token horizons.
5 Conclusions and future work
We present studies on several aspects of MXFP8-formats and their usage in pre-training models, e.g. scale computation, data format of specific tensors, numerical aspects of quantization, model layers that could be quantized for maximizing training throughput. We discuss recipes for pre-training LLMs with MXFP8-formats. The MXFP8 pre-training recipe is implemented in Transformer Engine [30] and the conversion numerics in cuDNN and cuBLAS libraries.
As part of our future work, we plan to extend our study to post-training phases (e.g. phases that perform RL based policy-optimizations and its variants [31] as well as supervised-finetuning stages). Additionally, from a hardware systems performance perspective, since tensors have to be quantized along two separate axes with MX-formats, the training framework needs to store two copies for each tensor (we discussed this in section 3.2). We plan to study scaling methods (e.g. 2D scaling instead of 1D) that would lower the storage overhead.
Acknowledgments
We thank members of the ADLR/PSX team (Sweta P., Mikail K., Ben L.) for helping with the draft revisions as well as Mohammad S., Carlo D.M., Michael A., Eric C. and Bryan C. with valuable feedback and discussions. We also thank PM for guidance throughout this work.
Appendix A Appendix
A.1 UE8M0 rounding
Computation described in Algorithm 1 is a simplification. The simplification comes from storing the output of $log_{2}$(float x) in FP32. $log_{2}$ function in device internally performs a round-to-nearest of the resulting return value and thus the result of ceil($log_{2}$(x)) can be different if the output of $log_{2}$ is not stored in a sufficiently large data-type. Hence, for emulation purposes we work directly with the bit representation of the ratio of amax and destmax. We, next, describe the computation flow.
Background: As a reminder, the quantization process from 32 high-precision values, $V_{i}$, to quantized values, $Q_{i}$; $1\leq i\leq 32$, is given by: $Q_{i}$ = Quantize_to_fp8$(V_{i}/2^{X})$. $2^{X}$ is the scale factor; $X$ is stored in an unsigned 8bit integer container in memory and interpreted as $2^{X}$ by the hardware. This scale factor decodes $Q_{i}$ back to $V_{i}$ (with quantization loss).
Value of scale factor ($2^{X}$) = float_to_8bits(amax/destmax), where amax is absolute maximum in input/source block (of 32 elements) and destmax is the largest positive number in the destination (MX) number system. float_to_8bits converts a floating-point number to a power-of-two number.
A float (FP32) number can be represented in IEEE convention as $2^{E}$$\times$ 1.mantissa (normal) or $2^{-126}$$\times$ 0.mantissa (denormal). E can lie between -127 to 127 (or 0 to 254 with the exponent bias) and can be represented in the 8-bit container for scale factor. -126 (or 1 with the exponent bias) is also representable by 8-bit container. mantissa lies between [0,1). So, the question is: should mantissa bits be rounded-up, rounded-down, round-to-nearest, discarded, etc. to create a power-of-two number? We find round-up to be the best choice for pre-training with MX-formats.
Rounding: For float_to_8bits(), the recommended order of computation is:
  1. 1.
    Compute the decoding scale as: decode_scale = block_amax/destmax
  2. 2.
    if decode_scale is below $2^{-127}$, then set it to $2^{-127}$ (which is the smallest value representable in UE8M0)
  3. 3.
    For all other values that are not powers of 2, round-up to the closest representable UE8M0 value.
By construction amax/destmax never exceeds $2^{127}$ (which is the largest value representable in UE8M0) with FP8, FP6 or FP4 formats. The above computations are done in the bit-space in emulation.
A.2 MX-format conversion
Section 2 relies on the standard MX-format conversion algorithm defined in [2], but for completeness we show it here given a shared scaling exponent $X$ as computed in 3.1.
Quantizing FP32 values to MX type: Once X is computed, $V_{i}/2^{X}$ is computed and the resulting value is quantized to a FP8-representable number (Quantize_to_fp8()). Round-to-nearest-ties-to-even (RN) rounding is used during this quantization step. The conversion process is saturating, i.e. if after rounding the resulting value exceeds FP8 max or is less than FP8 min value, then the result is clamped to respective max or min value.
Quantization operations add computation overhead — Blackwell has hardware support for rounding the scale (using our proposed method) and quantizing values to lower this overhead.
A.3 Why is special hardware needed for MX-formats?
Computing the matrix-product of two tensors involves performing dot-products between sub-vectors of the two tensors. Therefore, scaling factors need to be processed once per group of values that share the scale. Since MX-formats have fine-grained scaling, scale factors are processed once after each block’s dot-product is computed, thus, many times per tensor-wide dot-product. This is expensive to do in software, so hardware needs to add support for accelerating tensor operations involving MX-formats (e.g. Blackwell).
A.4 MXFP8 pre-training for a mixture-of-experts model
Section 3.3 discusses empirical data that shows MXFP8 matches BF16 accuracy (both training loss as well as downstream task accuracy). Transformer based mixture-of-experts (MoE) models are popular in literature. Fig. 6 shows that MXFP8 pre-training also matches BF16 pre-training loss curve for a MoE setup that we experimented with. The MoE model has 16 billion total parameters and $\sim$2.5 billion active parameters and we train the model on 1 trillion tokens. We follow the same guidelines discussed in section 3.2 for pre-training the MoE model. The pre-training phase uses a WSD [32] learning rate schedule. The final loss of the MXFP8-trained MoE model is within 0.1% of BF16-training.
Figure 6: MXFP8 versus BF16 for a MoE model
A.5 Model configurations
We conduct numerical experiments on LLM pre-training with variants of Nemotron-4 [11] models. Training and model details are described below. The 1T and 300B tokens dataset are a subset of the 17T data set discussed in [33]. Table 2 details the parameters for the various models that were used.
Table 2: Configuration for the LLM models.
Model Layers Hidden Attention Sequence Batch Initial Final
Size Heads Length Size LR LR
843M 24 1024 16 4096 256 2.5e-4 2.5e-7
2B 24 2048 16 4096 256 2e-4 2e-7
8B 32 4096 32 4096 1024 3e-4 3e-7